OpenStudy (hackergod):

Some investments in the stock market have earned 10% annually. At this rate , earnings can be found using the formua A=P(1.10)^n, where A is the total value of the investment, P is the initial value of the investment, and n is the number of years the money is invested. If 1,500 is invested in the stock marked at this annual rate of return, what is the expected total value after 18 years? A. 29,700.00 B. 28,050.00 C. 8,339.88 D. 7,581.71

OpenStudy (studygurl14):

All you do is use the formula given using the information you know, to find the answer

OpenStudy (hackergod):

What would it be? I will give you a medal and become a fan

OpenStudy (studygurl14):

A = P(1.10)^n A = total of the investment P = initial value = 1500 n = # of years invested = 18 Plug in values into equation A = 1500(1.10)^18 Can you do the rest? All you have to do is solve for A

OpenStudy (hackergod):

Would it be A

OpenStudy (studygurl14):

Nope. C